# Single-shot volumetric fluorescence imaging with neural fields

### v2

@@ Abstract
Single-shot volumetric fluorescence (SVF) imaging offers a significant advantage over traditional imaging methods that require scanning across multiple axial planes as it can capture biological processes with high temporal resolution across a large field of view. The key challenges in SVF imaging include requiring sparsity constraints to meet the multiplexing requirements of compressed sensing, eliminating depth ambiguity in the reconstruction, and maintaining high resolution across a large field of view. In this paper, we introduce the QuadraPol point spread function (PSF) combined with neural fields, a novel approach for SVF imaging. This method utilizes a custom polarizer at the back focal plane and a polarization camera to detect fluorescence, effectively encoding the 3D scene within a compact PSF without depth ambiguity. Additionally, we propose a reconstruction algorithm based on the neural fields technique that provides improved reconstruction quality and addresses the inaccuracies of phase retrieval methods used to correct imaging system aberrations. This algorithm combines the accuracy of experimental PSFs with the long depth of field of computationally generated retrieved PSFs. QuadraPol PSF, combined with neural fields, significantly reduces the acquisition time of a conventional fluorescence microscope by approximately 20 times and captures a 100 mm3 cubic volume in one shot. We validate the effectiveness of both our hardware and algorithm through all-in-focus imaging of bacterial colonies on sand surfaces and visualization of plant root morphology. Our approach offers a powerful tool for advancing biological research and ecological studies.
@@ 1 Introduction
Fluorescence imaging is an indispensable tool in biological research as it enables real-time observation of live organisms due to its high sensitivity, biological specificity, and non-invasive nature. However, conventional 2D fluorescence microscopy cannot capture the complete 3D structure of biological samples. To address this, techniques such as confocal and light sheet fluorescence microscopy have been developed and are widely used. However, these methods require scanning multiple axial planes, significantly increasing the acquisition time and limiting the spatial-temporal throughput. The extensive scanning time is impractical for numerous applications where imaging large fields of view (FOVs) is required, such as when performing lab studies of the complex biological processes associated with the rhizosphere [1] - the media-root interfaces of plants.
@@
While these approaches are effective, they also come with associated limitations. A significant challenge with these systems is that their point spread functions (PSFs) are not laterally shift-invariant, which necessitates extensive calibration to define the PSF accurately and restricts measurement to the pre-calibrated volume. Further, the shift-variant nature of the PSF complicates the image analysis – reconstructions typically require optimization algorithms that impose sparsity constraints. Another drawback of using coded masks or microlens arrays is that the large size decreases the peak signal level of the PSF by spreading photons over a larger area. Additionally, photons from various axial and lateral positions overlapping on the sensor further degrade the signal-to-noise ratio (SNR). To address some of these issues while improving resolution and depth of field (DOF), Miniscope3D [11] and Fourier DiffuserScope [12] replace the tube lens in a standard fluorescence microscope with an optimized phase mask, but they still suffer from some of the limitations associated with non-compact PSFs.
@@
PSF engineering is an alternative approach that modulates the fluorescence at the back focal plane (BFP) of the objective lens in the imaging system to encode the axial position of the emitters within compact PSFs. Numerous 3D PSFs have been proposed over the past two decades [13, 14], including the astigmatic PSF [15], which varies in elongation direction and magnitude with defocus; the corkscrew [16] and double-helix [17] PSFs, which feature revolving spots around the emitter; and the tetrapod [18] and pixOL [19] PSFs, which are optimized for maximizing the Fisher information. Another group of PSFs, including the bisected [20], quadrated [21], tri-spot PSFs [22], and the multi-view reflector (MVR) microscope [23], images subapertures off the pupil center to induce lateral displacements in the focal point; the displacement direction depends on the position of the subaperture relative to the center of the BFP, and the displacement amount is proportional to the defocus.
@@
While these engineered PSFs have proven effective primarily in single molecule localization microscopy [24, 25, 26, 27] where isolated point sources are imaged, reconstructing more complex geometries from these PSFs remains challenging. Unlike methods like Miniscope3D, compact PSFs usually cannot satisfy the multiplexing requirement of compressed sensing. Therefore, recovering the 3D object using a single image often results in ambiguities in depth measurements [28, 29]. Although using multiple images with engineered PSFs has the ability to substantially reduce these ambiguities, such as in the Fourier light field microscope [30, 31] and complex-field and fluorescence microscopy using the aperture scanning technique (CFAST) [29], these methods require either capturing multiple images sequentially at the expense of temporal resolution, or utilizing different areas of the detector to perform spatial multiplexing, thus sacrificing the FOV. A recent approach, the polarized spiral PSF [28], integrates polarizers with a double-helix phase mask and employs orthogonally polarized detection channels from a polarization camera to achieve single-shot 3D imaging without sacrificing either the temporal resolution or the FOV. However, it has not completely eliminated the ambiguity problem.
Another critical aspect in SVF imaging is the reconstruction algorithm; a robust reconstruction algorithm is essential for accurately and precisely reconstructing 3D scenes from 2D measurements captured with engineered PSFs. The Richardson-Lucy (RL) algorithm [30, 32, 33] is broadly used due to its effectiveness in recovering 3D structures. However, the reconstruction quality using this algorithm heavily depends on the accuracy of the forward model, particularly the PSF used. While experimentally calibrated PSFs can account for imperfections in the imaging system and improve the reconstruction quality, they can be inaccurate or unavailable due to system limitations or discrepancies in the forward imaging model. To address this limitation, we adopt the concept of neural fields in our work.
@@
Neural fields have become a prominent and prevalent technique for 3D scene representations and graphics rendering [34, 35]. Its concept focuses on the mapping between spatial coordinates and image intensity values using a compact multi-layer perceptron (MLP) model and optionally learnable positional embeddings. This approach has recently been applied to improve microscopic systems. For example, it has been implemented to represent continuous 3D refractive index in intensity diffraction tomography and to generate continuous images across the axial dimension [36], but it requires extensive optimization time on high-performance graphics processing units (GPU). Further developments reduced the computational resources by creating an additional hash encoding layer at the input of the neural network for 2D microscopic imaging systems [37]. More adaptations of the neural fields approach include a 2D version for lensless imaging phase retrieval [38] and the modeling of space-time dynamics with 4D neural fields in imaging through scattering [39], computational adaptive optics [40], and structured illumination microscopy [41]. Advancements continued with the exploitation of redundancies in Fourier ptychographic microscopes to augment the MLP model with a compact learnable feature space, thereby speeding up the image stack reconstructions and reducing data storage burdens [42]. Most recently, the integration of neural fields with diffusion models has been proposed to aid volumetric reconstruction [43].
@@ 2.1 Imaging system
Figure 1: (a) Schematic of single-shot volumetric fluorescence imaging using the QuadraPol PSF. OL, objective lens; TL, tube lens; QWP, quarter waveplate; DM, dichroic mirror; BF, bandpass filter. A 4-polarization custom polarizer (4-Pol) is positioned at the back focal plane (BFP) of the imaging system to modulate the emission light, with a polarization camera (PolCam) capturing the modulated fluorescence. The transmission axes of the polarizer and PolCam are 0°, 45°, 90°, and 135°. (b) Assembling the custom polarizer by aligning two coverslips and four laser-cut polymer polarizers between 3D-printed holders. (c) A representative image captured by the polarization camera, visualized using (i) raw pixel readouts, (ii) a polarization image color-coded in the HSV scheme (AoLP as hue, DoLP as saturation, and intensity as value), and (iii) four separate images for each polarization channel. Scale bar: 50 µm.
@@ 2.2 Point spread function design and experimental calibration
The design of the QuadraPol PSF is inspired by the quadrated PSF [21], the MVR microscope [23], and CFAST [29]. These techniques divide the pupil into four sections. Specifically, the quadrated phase mask applies different phase ramps to each segment, generating a four-spot PSF in a single imaging channel. In contrast, the MVR and CFAST systems utilize reflective mirrors and light-blocking apertures, respectively, to create distinct image channels for each pupil segment. A key characteristic of these techniques is the encoding of the axial position of the emitter through the different lateral displacements of the focused spots, each associated with one of the four equal segments of the pupil. However, these approaches come with limitations. The single-image capture with the quadrated PSF can lead to ambiguities in 3D reconstruction, while the MVR and CFAST setups sacrifice either FOV or temporal resolution. The QuadraPol PSF distinguishes itself by modulating each pupil section using a polarizer with a unique transmission axis, allowing the polarization camera to simultaneously capture four imaging channels without losing FOV or temporal resolution.
@@
Given the imaging system’s low NA, the theoretical PSF (Figure 2(a)) for the axial position $z$ in each polarization channel ($\mathrm{pol}\in\{0^{\circ},45^{\circ},90^{\circ},135^{\circ}\}$) can be described by (see Supplementary Section 1 for more details) [46, 47]
$\mathrm{PSF}_{z,\mathrm{pol}}=\sum_{i\in[p,s]}\mathcal{F}\left\{A_{\mathrm{pol},i}(u,v)\exp\left\{j\left[kz\sqrt{1-u^{2}-v^{2}}+P(u,v)\right]\right\}\right\}^{2},$ (1)
@@
where $k=2\pi/\lambda$ represents the wave number, and $(u,v)$ are the coordinates at the BFP. The pupil phase is consistent across all polarization channels; a uniform phase, $P(u,v)=0$, is used for calculating the unaberrated theoretical PSF. The amplitude modulation patterns for $p$- and $s$-polarizations are shown in Figure 2(a). These patterns assign values of 1, 0.5, and 0 for amplitude mask with $p$-polarization, and 0, $\pm$0.5, and 0 for $s$-polarization, corresponding to the angles between the transmission axis of the polarizers and detection channel at 0°, 45°, and 90°, respectively. Note that due to cross-talk between polarization channels separated by 45°, the PSF in each channel exhibits a primary spot and two weaker side spots, rather than a singular focal point (Supplementary Figure S1).
Figure 2: Amplitude and phase of the pupil and PSFs for at different heights. (a) Theoretical PSFs without aberration, (b) simulated PSFs using the retrieved phase, and (c) experimental PSFs. Scale bar: 2 mm for the pupil images, and 0.2 mm for the PSF images.
@@
Given the presence of aberrations in our imaging system, primarily introduced by the custom-made polarizer, we use the vectorial implementation of phase retrieval [48] to compensate for these imperfections. This approach enables us to refine a series of retrieved PSFs (Figure 2(b)) that more closely match those obtained experimentally (Figure 2(c)) compared with theoretically obtained ones (Figure 2(a)). These experimental PSFs are generated by capturing images from fluorescent beads (Thermo Fisher F8858) that were axially scanned over a 8 mm range with a 0.1 mm step size. The retrieved phase of the pupil is shown in Figure 2(b), while we assume that the pupil amplitude remains the same as those in Figure 2(a).
@@ 2.3 Reconstruction algorithm
The deconvolution using experimental PSFs, acquired from the calibration experiment at discrete $z$ planes, can produce accurate reconstruction results within a certain depth range. However, rhizosphere-asscociated samples, such as plant roots, can extend to a large depth range across different FOVs. To extend the depth range for reconstruction, the naive solution would be to obtain the PSFs at larger depths. Unfortunately, with the extended depth range, the PSF will expand to a larger size, resulting in low SNRs (Figure 2(c)). An alternative method is to implement the retrieved PSFs. Because they are generated from aberrated phase terms and ideal amplitude masks, they suit the needs at arbitrary $z$ planes with continuous sampling and can extend to a farther depth range. Nevertheless, the retrieved PSF cannot reconstruct high-quality image volumes as the experimental PSF via deconvolution. The discrepancy lies in the non-ideal pupil amplitude mask and inaccurate pupil phase retrieval.
@@
To address this dilemma, we customized and designed the neural fields to enhance the performance of our system. Given the deconvolved image volumes from experimental and retrieved PSFs, we started with the image volume merged from these two deconvolution results (Figure 3(a)) and began a two-stage optimization process. The first stage was to initialize our neural field with the merged image volume. The neural field learns a compact learnable feature space consisting of a 3D feature tensor ($M$) and a feature tensor ($u$) [42], and it also learns an MLP network including two nonlinear layers with ReLU activation functions and one linear layer. Each layer of the MLP in our neural field has $Q$ neurons with an additional offset neuron. The number of neurons is designed to match the number of feature channels in the feature space. The feature space is created from the Hadamard product of a feature tensor $M$ with a size of $N/2\times N/2\times Q$ and a feature tensor $u$ with a size of $Z\times Q$, where $N$ is the number of pixels along $x$ or $y$ axis of the raw image, $Q$ is the number of feature channels, and $Z$ denotes the number of pre-defined $z$-coordinates. The detailed parameters and hyperparameters of the neural fields can be found in Supplementary Section 2. The optimization process seeks to find a mapping function ($\phi$) between feature tensor space and the image volume:
$\phi(\mathbf{x}):\mathbb{R}^{2}\cdot\mathbb{R}\mapsto\mathbb{R}^{3},\;\;\mathbb{C}(f(\mathbf{x}),\phi(\mathbf{x}))$ (3)
@@
where $\mathbf{x}=(x,y,z)$ represents the 3D coordinate system and $\mathbb{C}$ is a set of constraints bounded by the merged image volume $f(\mathbf{x})$. After the first stage, the neural field can render an image volume as an initial guess ($g$).
Figure 3: Framework of using neural fields to extend the quality and depth range of the imaging system. (a) The initial image volume is obtained from the deconvolution using experimental PSFs (Exp. PSFs) for central frames and retrieved PSFs (Ret. PSFs) for peripheral frames. (b) The merged image volume from (a) guides the initialization of the model with a compact lernable feature space and multilayer perceptron (MLP). After model initialization, the model is further optimized for the image volume. The estimated image volume goes through the forward model of the imaging system to generate the estimated measurements. These measurements are compared with the acquired measurements, and then to update the model weights and parameters. (c) Once the model is optimized, the parameters and weights are fixed. It can render an image stack with continuous sampling. The operator $\otimes$ denotes a convolution operation, and $\oplus$ indicates a summation operation along $z$-axis.
At the second stage, we can render our estimated measurements:
$I_{\mathrm{pol}}(x,y)=\sum_{z}|g_{z}(\mathbf{x})\otimes\mathrm{PSF}_{z,\mathrm{pol}}|^{2}$ (4)
@@
where $I_{\mathrm{pol}}(x,y)$ is the estimated measurements from the neural field rendered image volume, $g_{z}$ denotes the rendered image volume at one $z$-plane, and $\otimes$ is the convolution operator. The PSFs are reused from the experimental PSF and retrieved PSF for an extended depth range. As our second stage of optimization, we optimized the neural field to minimize the difference between the estimated measurements and the experimental intensity measurements (Figure 3(b)).
@@
After the two-stage optimization, the parameters and weights in the neural field are fixed. We can continuously sample the feature tensor ($u$) to render a denser image volume along the axial ($z$) axis, as shown in Figure 3(c).

### v3

@@ Abstract
Single-shot volumetric fluorescence (SVF) imaging offers a significant advantage over traditional imaging methods that require scanning across multiple axial planes as it can capture biological processes with high temporal resolution. The key challenges in SVF imaging include requiring sparsity constraints, eliminating depth ambiguity in the reconstruction, and maintaining high resolution across a large field of view. In this paper, we introduce the QuadraPol point spread function (PSF) combined with neural fields, a novel approach for SVF imaging. This method utilizes a custom polarizer at the back focal plane and a polarization camera to detect fluorescence, effectively encoding the 3D scene within a compact PSF without depth ambiguity. Additionally, we propose a reconstruction algorithm based on the neural fields technique that provides improved reconstruction quality compared to classical deconvolution methods. QuadraPol PSF, combined with neural fields, significantly reduces the acquisition time of a conventional fluorescence microscope by approximately 20 times and captures a 100 mm3 cubic volume in one shot. We validate the effectiveness of both our hardware and algorithm through all-in-focus imaging of bacterial colonies on sand surfaces and visualization of plant root morphology. Our approach offers a powerful tool for advancing biological research and ecological studies.
@@ 1 Introduction
Fluorescence imaging is an indispensable tool in biological research as it enables real-time observation of live organisms due to its high sensitivity, biological specificity, and non-invasive nature. However, conventional 2D fluorescence microscopy cannot capture the complete 3D structure of biological samples. To address this, techniques such as confocal and light sheet fluorescence microscopy have been developed and are widely used. Despite their advantages, these methods require scanning multiple axial planes, significantly increasing the acquisition time and limiting the spatial-temporal throughput. The extensive scanning time is impractical for numerous applications where imaging large fields of view (FOVs) is required, such as when performing lab studies of the complex biological processes associated with the rhizosphere [1] - the media-root interfaces of plants.
@@
While these approaches are effective, they also come with associated limitations. A significant challenge with these systems, with the exception of the Fourier light field microscope [11, 12], is that their point spread functions (PSFs) are not laterally shift-invariant, which necessitates extensive calibration to define the PSF accurately and restricts measurement to the pre-calibrated volume. Further, the shift-variant nature of the PSF complicates the image analysis – reconstructions typically require optimization algorithms that impose sparsity constraints. Another drawback of using coded masks or microlens arrays is that the large size decreases the peak signal level of the PSF by spreading photons over a larger area. Additionally, photons from various axial and lateral positions overlapping on the sensor further degrade the signal-to-noise ratio (SNR). To address some of these issues while improving resolution and depth of field (DOF), Miniscope3D [13] and Fourier DiffuserScope [14] replace the tube lens in a standard fluorescence microscope with an optimized phase mask, but they still suffer from some of the limitations associated with non-compact PSFs.
@@
PSF engineering is an alternative approach that modulates the fluorescence at the back focal plane (BFP) of the objective lens in the imaging system to encode the axial position of the emitters within compact PSFs. Numerous 3D PSFs have been proposed over the past two decades [15, 16], including the astigmatic PSF [17], which varies in elongation direction and magnitude with defocus; the corkscrew [18] and double-helix [19] PSFs, which feature revolving spots around the emitter; and the tetrapod [20] and pixOL [21] PSFs, which are optimized for maximizing the Fisher information. Another group of PSFs, including the bisected [22], quadrated [23], and tri-spot PSFs [24], images subapertures off the pupil center to induce lateral displacements in the focal point; the displacement direction depends on the position of the subaperture relative to the center of the BFP, and the displacement amount is proportional to the defocus.
@@
While these engineered PSFs have proven effective primarily in single molecule localization microscopy [25, 26, 27, 28] where isolated point sources are imaged, reconstructing more complex geometries from these PSFs remains challenging. Unlike methods like Miniscope3D, compact PSFs usually cannot satisfy the multiplexing requirement of compressed sensing. Therefore, recovering the 3D object using a single image often results in ambiguities in depth measurements [29, 30]. Although using multiple images with engineered PSFs has the ability to substantially reduce these ambiguities, such as in the Fourier light field microscope [11, 12] and complex-field and fluorescence microscopy using the aperture scanning technique (CFAST) [30], these methods require either capturing multiple images sequentially at the expense of temporal resolution or utilizing different areas of the detector to perform spatial multiplexing, thus sacrificing the FOV. A recent approach, the polarized spiral PSF [29], integrates polarizers with a double-helix phase mask and employs orthogonally polarized detection channels from a polarization camera to achieve single-shot 3D imaging without sacrificing either the temporal resolution or the FOV. However, it has not completely eliminated the ambiguity problem.
@@
Another critical aspect of SVF imaging is the reconstruction algorithm; a robust reconstruction algorithm is essential for accurately and precisely reconstructing 3D scenes from 2D measurements captured with engineered PSFs. The Richardson-Lucy (RL) algorithm [11, 31, 32] is broadly used due to its effectiveness in recovering 3D structures. However, its performance degrades under noisy conditions, particularly in fluorescence imaging where the signal is limited. Given these limitations, it is worth exploring the use of neural fields in SVF. The topic of neural field is a recent prominent and prevalent technique for 3D scene representations and graphics rendering [33, 34]. Neural field techniques perform mapping between spatial coordinates and image intensity values using a compact multi-layer perceptron (MLP) model and optionally learnable positional embeddings. This approach has recently been applied to improve microscopic systems. For example, it has been implemented to represent continuous 3D refractive index in intensity diffraction tomography and to generate continuous images across the axial dimension [35], but it requires extensive optimization time on high-performance graphics processing units (GPU). Further developments reduced the computational resources by creating an additional hash encoding layer at the input of the neural network for 2D microscopic imaging systems [36]. More adaptations of the neural fields approach include a 2D version for lensless imaging phase retrieval [37] and the modeling of space-time dynamics with 4D neural fields in imaging through scattering [38], computational adaptive optics [39], and structured illumination microscopy [40]. Advancements continued with the exploitation of redundancies in Fourier ptychographic microscopes to augment the MLP model with a compact learnable feature space, thereby speeding up the image stack reconstructions and reducing data storage burdens [41]. Most recently, the integration of neural fields with diffusion models has been proposed to aid volumetric reconstruction [42].
Figure 1: (a) Schematic of single-shot volumetric fluorescence imaging using the QuadraPol PSF. OL, objective lens; TL, tube lens; QWP, quarter waveplate; DM, dichroic mirror; BF, bandpass filter. A 4-polarization custom polarizer (4-Pol) is positioned at the back focal plane (BFP) of the imaging system to modulate the emission light, with a polarization camera (PolCam) capturing the modulated fluorescence. The transmission axes of the polarizer and PolCam are 0°, 45°, 90°, and 135°. (b) Assembling the custom polarizer by aligning two coverslips and four laser-cut polymer polarizers between 3D-printed holders. (c) A representative image of a point source captured by the polarization camera, visualized using (i) raw pixel readouts, (ii) a polarization image color-coded in the HSV scheme (AoLP as hue, DoLP as saturation, and intensity as value), and (iii) four separate images for each polarization channel. Scale bar: 50 µm.
@@ 2.1 Imaging system
Figure 2: Amplitude and phase of the pupil and PSFs for at different heights. (a) Theoretical PSFs without aberration, (b) simulated PSFs using the retrieved phase, and (c) experimental PSFs. Question marks indicate that the phase and amplitude for the experimental PSF are not accessible. Scale bar: 2 mm for the pupil images, and 0.2 mm for the PSF images.
@@ 2.2 Point spread function design and experimental calibration
The design of the QuadraPol PSF is inspired by the quadrated PSF [23], the multi-view reflector (MVR) microscope [45], and CFAST [30]. These techniques divide the pupil into four sections. Specifically, the quadrated phase mask applies different phase ramps to each segment, generating a four-spot PSF in a single imaging channel. In contrast, the MVR and CFAST systems utilize reflective mirrors and light-blocking apertures, respectively, to create distinct image channels for each pupil segment. A key characteristic of these techniques is the encoding of the axial position of the emitter through the different lateral displacements of the focused spots, each associated with one of the four equal segments of the pupil. However, these approaches come with limitations. The single-image capture with the quadrated PSF can lead to ambiguities in 3D reconstruction, while the MVR and CFAST setups sacrifice either FOV or temporal resolution. The QuadraPol PSF distinguishes itself by modulating each pupil section using a polarizer with a unique transmission axis, allowing the polarization camera to simultaneously capture four imaging channels without losing FOV or temporal resolution.
Figure 3: Framework of using neural fields to extend the quality and depth range of the imaging system. (a) The RL-deconvolved image volume guides the initialization of the model with a compact learnable feature space and multilayer perceptron (MLP). After model initialization, the model is further optimized for the image volume. The estimated image volume goes through the forward model of the imaging system to generate the estimated measurements. These measurements are compared with the acquired measurements, and then to update the model weights and parameters. (b) Once the model is optimized, the parameters and weights are fixed. It can render an image stack with continuous sampling. The operator $\otimes$ denotes a convolution operation, and $\oplus$ indicates a summation operation along $z$-axis.
@@
Given the imaging system’s low NA, the theoretical PSF (Figure 2(a)) for the axial position $z$ in each polarization channel ($\mathrm{pol}\in\{0^{\circ},45^{\circ},90^{\circ},135^{\circ}\}$) can be described by (see Supplementary Note S1 for more details) [46, 47]
$\mathrm{PSF}_{z,\mathrm{pol}}=\sum_{i\in[p,s]}\mathcal{F}\{A_{\mathrm{pol},i}(u,v)\exp\{j[kz$ $\sqrt{1-u^{2}-v^{2}}$
$\hskip 12.0pt+P(u,v)\big{]}\big{\}}\big{\}}^{2},$ (1)
@@
where $k=2\pi/\lambda$ represents the wave number, and $(u,v)$ are the coordinates at the BFP. The pupil phase is consistent across all polarization channels; a uniform phase, $P(u,v)=0$, is used for calculating the unaberrated theoretical PSF. The amplitude modulation patterns for $p$- and $s$-polarizations are shown in Figure 2(a). These patterns assign values of 1, 0.5, and 0 for amplitude mask with $p$-polarization, and 0, $\pm$0.5, and 0 for $s$-polarization, corresponding to the angles between the transmission axis of the polarizers and detection channel at 0°, 45°, and 90°, respectively. Note that due to cross-talk between polarization channels separated by 45°, the PSF in each channel exhibits a primary spot and two weaker side spots, rather than a singular focal point (Supplementary Figure S1). These images exhibit apparent lateral shifts in different directions across different polarization channels as the object defocuses; this is analogous to viewing the same object from four unique perspectives, which fundamentally enables 3D reconstruction.
@@
Given the presence of aberrations in our imaging system, primarily introduced by the custom-made polarizer, we use the vectorial implementation of phase retrieval [48] to compensate for these imperfections. This approach enables us to refine a series of retrieved PSFs (Figure 2(b)) that more closely match those obtained experimentally (Figure 2(c)) compared with theoretically obtained ones (Figure 2(a)). These experimental PSFs are generated by capturing images from fluorescent beads (Thermo Fisher F8858) that were axially scanned over an 8 mm range with a 0.1 mm step size. The retrieved phase of the pupil is shown in Figure 2(b), while we assume that the pupil amplitude remains the same as those in Figure 2(a). Both experimental and retrieved PSFs are implemented in the image reconstruction in Section 3.4.
@@ 2.3 Reconstruction algorithm
One limitation of the RL deconvolution is its performance under noisy situations. Inspired by the forward imaging model, which captures four independent perspectives of the 3D scene, we customized and designed the neural fields to enhance the performance of our system. The algorithm is composed of a two-stage optimization process. The first stage was to initialize our neural fields with the RL-deconvolved image volume. The neural field learns a compact learnable feature space consisting of a 3D feature tensor ($M$) and a feature tensor ($u$) [41], and it also learns an MLP network including two nonlinear layers with ReLU activation functions and one linear layer. Each layer of the MLP in our neural field has $Q$ neurons with an additional offset neuron. The number of neurons is designed to match the number of feature channels in the feature space. The feature space is created from the Hadamard product of a feature tensor $M$ with a size of $N/2\times N/2\times Q$ and a feature tensor $u$ with a size of $Z\times Q$, where $N$ is the number of pixels along $x$ or $y$ axis of the raw image, $Q$ is the number of feature channels, and $Z$ denotes the number of pre-defined $z$-coordinates. The detailed parameters and hyperparameters of the neural fields can be found in Supplementary Note S2 [49, 50]. The optimization process seeks to find a mapping function ($\phi$) between feature tensor space and the image volume
$\phi(\mathbf{x}):\mathbb{R}^{2}\cdot\mathbb{R}\mapsto\mathbb{R}^{3},\;\;\mathbb{C}(f(\mathbf{x}),\phi(\mathbf{x})),$ (3)
@@
where $\mathbf{x}=(x,y,z)$ represents the 3D coordinate system and $\mathbb{C}$ is a set of constraints bounded by the image volume $f(\mathbf{x})$. This initialization process can significantly shorten the rendering time. After the first stage, the neural field can render an image volume as an initial guess ($g$).
At the second stage, we can render our estimated measurements
$I_{\mathrm{pol}}(x,y)=\sum_{z}|g_{z}(\mathbf{x})\otimes\mathrm{PSF}_{z,\mathrm{pol}}|^{2},$ (4)
@@
where $I_{\mathrm{pol}}(x,y)$ is the estimated measurements from the neural field rendered image volume, $g_{z}$ denotes the rendered image volume at one $z$-plane, and $\otimes$ is the convolution operator. At our second stage of optimization, we optimized the neural field to minimize the difference between the estimated measurements and the experimental intensity measurements (Figure 3(a)).
Figure 4: Performance evaluation of the QuadraPol PSF using simulated data. (a) Lateral and (b) axial resolutions determined by the Rayleigh criterion as functions of axial position and signal level. Images show representative data with Poisson shot noise and reconstruction cross-sections using RL deconvolution and neural fields. Shaded areas represents the diffraction ($\lambda/2\mathrm{NA}$) and sampling limits. Scale bar: 20 µm in $xy$ view and 200 µm in $xz$ view. (c-e) Reconstruction results for various line structures; insets show simulated images. Parameters for simulating the QuadraPol PSF ($\mathrm{NA}=0.52$, magnification $M=-5.2$, camera pixel size $2.2$ µm) are adjusted to match those reported for Miniscope3D [13]. Scale bar: 100 µm; color bar: height in µm. (f,g) Object and images using the polarized spiral (PS2F) and QuadraPol PSFs with (f) vertical and (g) horizontal lines at $z=\pm 0.35$ mm. Scale bar: 50 µm; color bar: AoLP. Insets show the PSFs. (h) Comparison between the QuadraPol PSF and other SVF techniques.
@@
After the two-stage optimization, the parameters and weights in the neural field are fixed. We can continuously sample the feature tensor ($u$) to render a denser image volume along the axial ($z$) axis, as shown in Figure 3(b). Throughout the whole reconstruction process, no pre-training or training datasets are needed, and the algorithm is only supervised by a single capture. The average optimization time over 196 FOVs in the plant root experiment (see Section 3.4 for more details) on an Nvidia A100-80GB GPU device is $\sim$22 seconds.